# A Class of Countably Covered Two-Dimensional Sofic Shifts

### v1

@@ 4 Construction toolbox
Figure 5: A configuration over $A_{\mathrm{probe}}$, superimposed over a configuration of $X_{\mathrm{grid}}$. Each vertical segment is labeled with the type of symbol it carries; $a_{i}$ stands for $A\times\{i\}$.
A weaker version of the following result was stated explicitly in [10].
@@
Let $g:\mathbb{N}\to\mathbb{N}$ be a partial computable function. Then there exists a sequence $(C_{n},T_{n},a_{n},b_{n},t_{n})_{n\in\mathbb{N}}$, computable in polynomial time, with the following properties for all $n\in\mathbb{N}$.
@@
    $a_{n},b_{n}\in C_{n}$ are edge colors and $t_{n}\in T_{n}$ a tile such that for some $m\geq g(n)$, $T_{n}$ tiles an $m\times m$ rectangle whose entire north border is colored with $a_{n}$, west border with $b_{n}$, and the tile at the southeast corner is $t_{n}$. Furthermore, $T_{n}$ does not tile any such rectangle for $m<g(n)$.
@@
For $n\in\mathbb{N}$, let $M$ be a Turing machine that counts to some number $k\geq g(n)$, enters a special state $q$ that is not used otherwise, counts to $k$ again and then halts. We construct the machine in such a way that it has $\lceil\log n\rceil+K$ states for some constant $K$, by having it first write $n$ in binary on its tape and then perform a uniform algorithm with input $n$.
In [5], Kari constructs a NW-deterministic version of the aperiodic Robinson tile set [8] and embeds an arbitrary Turing machine in it. Its tilings consist of nested square patterns, called $(2^{n}-1)$-squares, each of which hosts a simulation of the embedded machine for approximately $n/2$ steps. If the machine halts in these $n$ steps, the pattern cannot be tiled correctly.
Denote by $R_{n}$ this tile set with $M_{n}$ being the embedded machine. It can tile a rectangle of side length $g(n)$ whose entire north border is colored by one color $a_{n}$ – the one that occurs at the north border of every southeast-facing $(2^{n}-1)$-square – and west border by the color $b_{n}$ found on the west border of these squares. The tile that simulates the special state $q$ is chosen as $t_{n}$; we can build $M_{n}$ and the simulation so that it always lies on the diagonal of a $(2^{n}-1)$-square. Then item 2 is satisfied. Since the machine $M_{n}$ eventually halts, $R_{n}$ does not tile the plane. Finally, the number of colors in $R_{n}$ grows linearly in the number of states and transitions of $M_{n}$, and we can make the latter grow as $O(\log\log n)$ by replacing $g$ with $g\circ\exp$ and using $M_{\lceil\log n\rceil}$ in place of $M_{n}$. Note that the sequence is allowed to have repeated values. ∎
The following lemma is a key ingredient of our construction. It allows us to take a small grid and use it to enforce the existence of a large but still finite grid on another layer.
@@
The tile set $S_{n}$ does not admit tilings without a finite grid, since $T_{n}$ does not tile the infinite plane. By compactness, the set of grid sizes it admits is finite. It does admit at least one tiling with a grid of some width $m\geq g(2^{n})$, but not of any width $m<g(2^{n})$. Also, the decorations inside each grid cell forms the same pattern in any tiling, since $T_{n}$ is NW-deterministic and the colors north and west borders of the grid cells are fixed. Thus the set of valid tilings over $S_{n}$ is countable.
@@
We describe the SFT $X_{g}$. Let $M$ be a counter machine with six named counters $b,m,t_{e},t_{n},t_{w},t_{s}$ that checks whether the quadruple $t=(t_{e},t_{n},t_{w},t_{s})\in\mathbb{N}^{4}$ is a tile of $S_{m}$, and that $b=1$ when $t$ corresponds to a tile with a horizontal grid line and $b=0$ otherwise. Let $h(m)\geq m$ be an upper bound for the number of computation steps required for this check for a fixed $m$ and any values of the other named counters. By Lemma 16 and the exponential slowdown in simulating a Turing machine by a counter machine, we can choose $h(m)=2^{m^{q}}$ for some $q\in\mathbb{N}$. We may also assume that the machine always takes the same number of steps for the check when $m$ is fixed. For example, we may have it perform the check for all possible values of the counters $t_{e}$, $t_{n}$, $t_{w}$ and $t_{s}$ that are below the maximum value $\log\log m$ and ignore all results except the one with the actual counter values; the total number of steps is still $2^{n^{O(1)}}$, so this does not violate the previous assumption. After $M$ has finished the check, it runs for at least $h(m+1)$ more steps (e.g. by simply counting to $h(m+1)$) and then halts.
@@
Similarly to Example 14, let $X\subset X_{\mathrm{grid}}\times X_{\mathrm{grid}}\times A_{g}^{\mathbb{Z}^{2}}$ be the SFT where each grid cell of the first component contains a simulated computation of $M$ on the third component, and adjacent grid cells synchronize the simulated counters so that each machine has the same value of $m$ and the colors of the simulated tiles match. We allow the simulation of $M$ to be cut at any point after the check but before it has halted. For each cross tile $t$ of the first component, we require that $t$ is matched with a horizontal line on the second component if and only if $t$ is adjacent to a simulated $b$-counter with value $1$ on the $A_{g}$-component. Also, each cross tile of the second component must be matched with a cross tile on the first component. For finitely many $n$ (including those for which the grid cells of $x_{\mathrm{grid}}(n)$ are too small to host a computation), we can enforce the contents of the second component by explicit rules. This concludes the definition of $X_{g}$. We call the three components of $X_{g}$ the small grid layer, the large grid layer and the simulation layer.
@@
We show that $X_{g}$ has the desired properties. Take any $n\geq 1$. Let $y$ and $z$ be such that $z^{\prime}=(x_{\mathrm{grid}}(n),y,z)\in X_{g}$. If $n$ is large enough, every grid cell of the small grid hosts a simulation of $M$ that computes a tile from the same set $S_{m}$ and enforces the adjacency rules to its neighbors. We must have $m\geq\log n$, since $M$ always runs for at least $m$ steps, which corresponds to at least $2^{m}$ rows in the tiling.
@@
By Lemma 16, every valid tiling over $S_{m}$ contains a grid of finite width. Hence the small grid layer contains regularly spaces cross tiles that must be matched with horizontal lines on the large grid layer $y$. Thus $y$ has a finite-width grid, that is, it is a shifted version of some $x_{\mathrm{grid}}(k)$ for $k\in\mathbb{N}$. Moreover, the simulated grid of $S_{m}$ has width at least $g(2^{m})\geq g(2^{\log n})=g(n)$, so that of $y$ has width at least $ng(n)>g(n)$ for large enough $n$ (and small values of $n$ are handled separately). As the two grids must be aligned, we have $y=\sigma^{\vec{v}}(x_{\mathrm{grid}}(k))$ for some $k\geq g(n)$ and $\vec{v}\in n\mathbb{Z}^{2}$. Since $x_{\mathrm{grid}}(n)$ is $n$-periodic both horizontally and vertically, $\vec{v}$ can be chosen arbitrarily. Thus there exists some set $K\subset\{g(n),g(n+1),\ldots\}$ such that the claimed equation holds.
@@
5 Proof of the main result
We now combine our puzzle pieces in order to form a proof of Theorem 6.
Proof of Theorem 6.
One direction follows from Theorem 3. For the converse, suppose we are given a nondecreasing computable function $f:\mathbb{N}\to\mathbb{N}$ and a shift space $X\subseteq G(f)$ that satisfies the countable cover conditions. Our goal is to construct a countable SFT cover for $X^{\dagger}$. In the first phase of the construction, we embed $X^{\dagger}$ into a countably covered sofic shift $Y$ that is contained in $G(g)^{\dagger}$ for some function $g:\mathbb{N}\to\mathbb{N}$ that majors $f$. After this, we will further restrict the SFT cover of $Y$, transforming it into a cover of $X$.
We begin to construct a countable SFT $Z$ and a projection $\pi:Z\to A^{\mathbb{Z}^{2}}$ such that $Y=\pi(Z)$ satisfies $G(f)^{\dagger}\subseteq Y\subseteq G(g)$ for some function $g$. This SFT consists of several layers, the first of which is the vertically constant $A$-layer $(A^{\dagger})^{\mathbb{Z}}$. We define $\pi$ as the projection to this layer.
The second layer is a copy of $X_{\mathrm{grid}}$, which we call the small grid layer. The third layer is the width restriction layer given by the tiles visible in Figure 6 and their obvious matching rules. Its thick horizontal lines must coincide with those of the small grid layer, and its vertical lines must coincide with the nonzero columns of the $A$-layer. It restricts the side length of the small grid layer to be at most the maximum distance between two nonzero columns of the $A$-layer. Note that the width restriction layer is completely determined by the previous layers in the case that there are at least two nonzero columns in the $A$-layer, and has countably many configurations otherwise.

### v2

@@ 4 Construction toolbox
Figure 5: A configuration over $A_{\mathrm{probe}}$, superimposed over a configuration of $X_{\mathrm{grid}}$. Each vertical segment is labeled with the type of symbol it carries (each tile of the segment carries the same symbol); $a_{i}$ stands for $A\times\{i\}$.
Consider a configuration $(x,y,z)\in H$, where $x\in X$ contains a grid of width $m\geq 1$ that simulates a configuration of tile set $T_{n}$ for some $n\in\mathbb{N}$, and $y\in(A^{\mathbb{Z}})^{\dagger}$ is the base alphabet layer. By rule 1, the horizontal and vertical lines of the probe implementation layer $z$ are aligned with the grid. Each grid cell $C$ contains a simulated copy of the machine $M$, which enters exactly one probe state $q_{b}$ during its computation. Here, $b\in B$ is the probe interface layer of the tile simulated by $C$. At the point where the zig-zag head of $C$ touches the west border of its computation cone in state $q_{b}$, rule 3 causes the neighboring tile on the west border of $C$, and hence the entire west border, to be colored with the symbol $b$.
By our assumption on $T_{n}$, the probe interface layer of each row of the simulated tiling contains at most one symbol from $A\times\{0,1\}$, and on its right some number of $1$-symbols, all other symbols coming from $\{0,2\}$. Thus, a given row of the grid contains at most one grid cell $C_{0}$ whose west border holds a value $(a,c)\in A\times\{0,1\}$. Cell $C_{0}$ is the one whose $A$-layer we are probing on this particular row. Thus, if the next $k\geq 0$ grid cells $C_{1},\ldots,C_{k}$ to the east of $C_{0}$ hold $1$-symbols on their west borders, we need to access the base alphabet layer of the $k$th column of $C_{0}$ from the left. We describe how this is achieved with the help of the probe implementation layer.
Suppose first that $k\geq 1$. Note from Figure 4 that in a crossing of grid lines (last 4 tiles on row 1), the north border of the tile holds a $1$-symbol if and only if the northwest part of the tile is light gray. Hence, a grid cell contains a gray region if and only if its east border holds a $1$-symbol. The gray region of each grid cell that contains one is a rectangle delimited by its south, west and east borders and a dashed horizontal line. If two neighboring grid cells contain gray regions, their heights differ by 1, the west one being higher (see tiles 4 and 5 on row 3 of Figure 4).
The east border of grid cell $C_{k}$ does not hold a $1$-symbol, so $C_{k}$ does not contain a gray region. Inductively, for each $i=0,\ldots,k-1$ the grid cell $C_{i}$ has a gray region of height kik-i. From the northwest corner of the gray region of $C_{0}$, which has height $k$, a diagonal signal is sent to the southeast (tile 6 on row 3 of Figure 4).
The diagonal signal carries the same symbol $(a,c)$ as the west border of $C_{0}$ (tiles 6 and 7 on row 3 of Figure 4). As it hits the south border of $C_{0}$ at distance $k$ from its southwest corner (tiles 6 and 7 on row 2 of Figure 4), rule 2 guarantees that the symbol on the base alphabet layer at that position is $a$. Also, since the diagonal signal cannot intersect another vertical grid line, we must have $k<m$. Because of the black disks at the borders of certain tiles, we have $c=1$ if and only if the diagonal signal hits the rightmost tile of the south border of $C_{0}$, if and only if k=m1k=m-1.
Suppose now that $k=0$, so that the east border of $C_{0}$ holds the value $2$ and $C_{0}$ does not contain a gray region. The southwest corner of $C_{0}$ must then be the tile at the top right of Figure 4 with $b_{0}\in A\times\{0\}$. By rule 2, the west border of $C_{0}$ holds the symbol $(a,0)$, where $a\in A$ is the symbol of the base alphabet layer on the same column.
We have now shown that for each row of a simulated tiling of $T_{n}$ whose probe interface layer contains a word of the form $(a,c)1^{k}2$ with $k\geq 0$, we must have $k<m$, $a$ must equal the symbol on the $A$-layer of the $k$th column of the grid cell simulating the symbol $(a,c)$, and $c=1$ if and only if k=m1k=m-1.
As for the cardinality of the new SFT, note first that if the grid layer contains a finite-width grid, then the probe implementation layer is completely determined by the other layers. Conversely, if there is no finite-width grid, then the probe implementation layer too contains no parallel thick lines, and a simple case analysis shows that the number of valid contents for this layer is countable when the other layers are fixed.
A weaker version of the following result is stated explicitly, and proved in less detail, in [12] as part of the proof of Theorem 1. The new properties are the tile $t_{n}$ and the bound $k_{n}=O(\log\log n)$. The tileset needs to be modified to obtain them.
@@
Let $g:\mathbb{N}\to\mathbb{N}$ be a computable function. Then there exists a sequence $(C_{n},T_{n},a_{n},b_{n},t_{n})_{n\in\mathbb{N}}$, computable in polynomial time, with the following properties for all $n\in\mathbb{N}$.
@@
    $a_{n},b_{n}\in C_{n}$ are edge colors and $t_{n}\in T_{n}$ a tile such that for some $m\geq g(n)$, $T_{n}$ tiles an $m\times m$ square whose entire north border is colored with $a_{n}$, west border with $b_{n}$, and the tile at the southeast corner is $t_{n}$. Furthermore, $T_{n}$ does not tile any such rectangle for $m<g(n)$.
@@
For $n\in\mathbb{N}$, let $M_{n}$ be a Turing machine that counts to $g(n)$, enters a special state $q$ that is not used otherwise, counts to $g(n)$ again and then halts. We construct the machine in such a way that it has $\lceil\log n\rceil+K$ states for some constant $K$, by having it first write $n$ in binary on its tape and then perform a uniform algorithm with input $n$. Let $t(n)\geq 2g(n)$ be the number of steps $M_{n}$ takes before halting.
In [6], Kari constructs a NW-deterministic version of the aperiodic Robinson tile set [10] and embeds an arbitrary Turing machine in it. Its tilings consist of nested square patterns, called $(2^{k}-1)$-squares, each of which hosts a simulation of the embedded machine for approximately $k/2$ steps. If the machine halts in these $k/2$ steps, the pattern cannot be tiled correctly.
Denote by $R_{n}$ this tile set with $M_{n}$ being the embedded machine. It can tile a southeast-facing $(2^{k}-1)$-square with $2^{k-1}\leq t(n)/2<2^{k}$. The north border of such a square always has the same color $a_{n}$, and its west border always has the same color $b_{n}$. The tile that simulates the special state $q$ is chosen as $t_{n}$; we can build $M_{n}$ and the simulation so that it always lies on the diagonal of a $(2^{n}-1)$-square. Then item 2 is satisfied with the $m\times m$-pattern that spans the northwest corner of the $(2^{k}-1)$-square and the simulated state $q$ (which takes at least $g(n)$ steps to reach, whence $m\geq g(n)$).
Since the machine $M_{n}$ eventually halts, $R_{n}$ does not tile the plane. Finally, the number of colors in $R_{n}$ grows linearly in the number of states and transitions of $M_{n}$, and we can make the latter grow as $O(\log\log n)$ by replacing $g$ with $g\circ\exp$ and using $M_{\lceil\log n\rceil}$ in place of $M_{n}$. Note that the sequence is allowed to have repeated values. ∎
The following lemma is a key ingredient of our construction. It allows us to take a small grid and use it to enforce the existence of a large but still finite grid on another layer. The idea is implicitly present in the proof of [12, Theorem 3], but used in an ad hoc manner and not formalized.
@@
The tile set $S_{n}$ does not admit tilings without a finite grid, since $T_{n}$ does not tile the infinite plane. By compactness, the set of grid sizes it admits is finite. It does admit at least one tiling with a grid of some width $m\geq g(2^{n})$, but not of any width $m<g(2^{n})$. Also, the decorations inside each grid cell form the same pattern in every tiling, since $T_{n}$ is NW-deterministic and the colors of the north and west borders of the grid cells are fixed. Thus the set of valid tilings over $S_{n}$ is countable.
Similarly to Example 17, we define $X_{g}\subset X_{\mathrm{grid}}\times X_{\mathrm{grid}}\times A_{g}^{\mathbb{Z}^{2}}$, where $A_{g}$ is an auxiliary alphabet, as an SFT where each grid cell of the first component contains a simulated computation of a counter machine $M$ on the $A_{g}$-component. We call the three components of $X_{g}$ the small grid layer, the large grid layer and the simulation layer.
The machine $M$ has six named counters $b,m,t_{e},t_{n},t_{w},t_{s}$. It nondeterministically chooses some value for each counter (with $b\in\{0,1\}$), and checks whether the quadruple $t=(t_{e},t_{n},t_{w},t_{s})\in\mathbb{N}^{4}$ is a tile of $S_{m}$. As in Example 17, adjacent small grid cells synchronize these simulated counters so that each machine has the same value of $m$ and the colors of the simulated tiles match, so that a configuration containing a finite-width small grid will simulate a configuration of some $S_{m}$.
@@
After this, the machine $M$ checks that $b=1$ when $t$ corresponds to a tile of $S_{m}$ with a horizontal grid line, and $b=0$ otherwise. Let $h(m)\geq m$ be an upper bound for the number of computation steps required thus far for a fixed $m$ and any values of the other named counters. By Lemma 19 and the exponential slowdown in simulating a Turing machine by a counter machine, we can choose $h(m)=2^{m^{q}}$ for some $q\in\mathbb{N}$. We may also assume that the machine always takes the same number of steps for the check when $m$ is fixed. For example, we may have it perform the check for all possible values of the counters $t_{e}$, $t_{n}$, $t_{w}$ and $t_{s}$ that are below the maximum value $\log\log m$ and ignore all results except the one with the actual counter values; the total number of steps is still $2^{n^{O(1)}}$, so this does not violate the previous assumption.
After $M$ has finished the check, it runs for at least $h(m+1)$ more steps, e.g. by simply computing $h(m+1)$ and then counting to $h(m+1)$. After this it finally halts. Note that the time required to compute $h(m+1)$ is irrelevant, since we only need $M$ to run for at least $h(m+1)$ steps but not indefinitely. We allow the simulation of $M$ to be cut by a horizontal small grid line at any point after it has finished its check (that is, after $h(m)$ computation steps) but before it has halted.
@@
For each cross tile $t$ of the small grid, we require that $t$ is matched with a horizontal line on the large grid if and only if the northeast neighbor of $t$ contains a simulated $b$-counter with value $1$ on its $A_{g}$-component. Also, each cross tile of the large grid must be matched with a cross tile on the small grid. For finitely many widths of the small grid (including those for which the small grid cells are too small to host a computation), we can enforce the width of the large grid by explicit rules. This concludes the definition of $X_{g}$.
@@
We show that $X_{g}$ has the desired properties. Take any $n\geq 1$. Let $y$ and $z$ be such that $z^{\prime}=(x_{\mathrm{grid}}(n),y,z)\in X_{g}$. If $n$ is large enough, every small grid cell (which has size $n\times n$) hosts a simulation of $M$ that computes a tile from the same set $S_{m}$ and enforces the adjacency rules to its neighbors. We must have $m\geq\log n$, since $M$ always runs for at least $m$ steps, which corresponds to at least $2^{m}$ rows in the tiling.
@@
By the second paragraph of this proof, the tiling of $S_{m}$ simulated on the small grid cells contains a grid of some finite width $w\geq g(2^{m})$. Hence the small grid layer contains regularly spaced small grid cells that simulate cross tiles. These must be matched with horizontal lines on the large grid layer $y$. Thus $y$ has a finite-width large grid, that is, it is a shifted version of some $x_{\mathrm{grid}}(k)$ for $k\in\mathbb{N}$. Moreover, the simulated grid of $S_{m}$ has width $w\geq g(2^{m})\geq g(2^{\log n})=g(n)$, so the large grid of $y$ has width at least $ng(n)>g(n)$ for large enough $n$ (and small values of $n$ are handled separately). As the small and large grids must be aligned, we have $y=\sigma^{\vec{v}}(x_{\mathrm{grid}}(k))$ for some $k\geq g(n)$ and $\vec{v}\in n\mathbb{Z}^{2}$. Since the small grid $x_{\mathrm{grid}}(n)$ is $n$-periodic both horizontally and vertically, $\vec{v}$ can be chosen arbitrarily. Thus there exists some set $K\subset\{g(n),g(n)+1,\ldots\}$ such that the claimed equation holds.
@@
5 Proof of Theorem 6
We now combine our puzzle pieces in order to form a proof of Theorem 6. One direction follows directly from Theorem 3.
For the converse, suppose we are given a nondecreasing computable function $f:\mathbb{N}\to\mathbb{N}$ and a shift space $X\subseteq G(f)$ that satisfies the countable cover conditions. Our goal is to construct a countable SFT cover for $X^{\dagger}$. In the first phase of the construction, we embed $X^{\dagger}$ into a countably covered sofic shift $Y$ that is contained in $G(g)^{\dagger}$ for some function $g:\mathbb{N}\to\mathbb{N}$ that dominates $f$. After this, we will further restrict the SFT cover of $Y$, transforming it into a cover of $X$.
Since the construction is rather complex and contains many interconnected layers, we list here all the relevant layers, their roles, and the place where they are defined. The reader may use this list as a quick reference.
  • Base alphabet layer: $(A^{\mathbb{Z}})^{\dagger}$, target of the projection. Defined in Section 5.1.
  • Small grid layer: its width $n$ is at most the minimum distance between nonzero columns of the base alphabet layer. Defined in Section 5.1.
  • Width restriction layer: implements the restriction on the small grid layer. Defined in Section 5.1.
  • Large grid layer: has width $m\geq f(n)+2$ (using Lemma 20). Define $g(n)$ as the maximum width of this layer. Defined in Section 5.1.
  • Counting layer: restricts the number of nonzero columns of the base alphabet layer to be at most $m$ (using Lemma 15). Defined in Section 5.1.
  • Simulation grid: has width at least $t(n)\cdot m$ (using Lemma 20), simulates a tiling over tile set $T_{n}$. Defined in Section 5.3.
  • Probe interface layer of $T_{n}$ and probe implementation layer: together allow the simulated tilings over $T_{n}$ to probe the base alphabet layer. Defined in Example 18 and Section 5.4.
  • Layer $L_{n,a,b}$ of $T_{n}$: responsible for checking the correctness of the base alphabet layer between nonzero columns $a$ and $b$. There is one for each $0\leq a<b\leq f(n)-1$. Defined in Section 5.4.
  • Computation sublayer of $L_{n,a,b}$: simulates a counter machine $C_{n}$ that controls the probing and correctness checks in $L_{n,a,b}$. Defined in Section 5.5.
  • Anchor sublayer of $L_{n,a,b}$: has the form $\cdots 001^{k}22\cdots$, where the $1$-symbols cover the area between nonzero columns $a$ and $b$ of the base alphabet layer. Defined in Section 5.5.
  • Small restart grid sublayer of $L_{n,a,b}$: has width $1\leq w\leq k$, where $k$ is given by the anchor sublayer. Defined in Section 5.7.
  • Large restart grid sublayer of $L_{n,a,b}$: has width at most $h^{\prime}(w)$ (using Lemma 20), where the simulated computation of $C_{n}$ fits on $h^{\prime}(k)$ rows. Periodically restarts the computation of $C_{n}$. Defined in Section 5.7.
5.1 Embedding $X^{\dagger}$ in a countably covered $G(g)^{\dagger}$
We begin to construct a countable SFT $Z$ and a projection $\pi:Z\to A^{\mathbb{Z}^{2}}$ such that $Y=\pi(Z)$ satisfies $G(f)^{\dagger}\subseteq Y\subseteq G(g)$ for some function $g$. This SFT consists of several layers, the first of which is the vertically constant $A$-layer $(A^{\mathbb{Z}})^{\dagger}$, called the base alphabet layer. We define $\pi$ as the projection to this layer.
The second layer is a copy of $X_{\mathrm{grid}}$, which we call the small grid layer. The third layer is the width restriction layer given by the tiles visible in Figure 6 and their obvious matching rules. Its thick horizontal lines must coincide with those of the small grid layer, and its vertical lines must coincide with the nonzero columns of the base alphabet layer. It restricts the width of the small grid layer to be at most the minimum distance between two nonzero columns of the base alphabet layer. In particular, if there are at least two nonzero columns, then the small grid layer will necessarily contain a grid of finite width (and the width can be any integer between $1$ and the minimum distance). Note that the width restriction layer is completely determined by the previous layers in the case that there are at least two nonzero columns in the base alphabet layer, and has countably many configurations otherwise.